OpenStudy (chrisneedshelp):

.Find the vertex of the parabola whose equation is y = x2 + 2x + 9. (-1, 8) (1, 12) (0, 9)

OpenStudy (anonymous):

The vertex of the parabola is the point where the axis and parabola intersect. Its x-coordinate \(x=-\dfrac{b}{2a}\) is its y-coordinate is found by substituting \(x=\frac{-b}{2a}\) in the parabola’s equation.

OpenStudy (anonymous):

So, compare your given eqn to the quadratic expression \(ax^2 + bx + c \) and you find \(a = 1, b=2, c=9\) x-coordinate of vertex = \(\dfrac{-2}{2*1} = -1\) y-coordinate of vertex = \(\small (-1)^2 + 2(-1) + 9 = 1 -2 + 9\rightarrow 8\) Vertex: (-1, 8)

OpenStudy (chrisneedshelp):

thank you again d3Xter! :)

OpenStudy (anonymous):

np :)